# The fate of entanglement

### v1

@@ Abstract
Quantum entanglement is a fundamentally non-local correlation between particles. In its simplest realisation, a measurement on one particle is affected by a prior measurement on its partner, irrespective of their separation. For multiple particles, purely collective types of entanglement exist but their detection, even theoretically, remains an outstanding open question. Here, we show that all forms of multi-party entanglement entirely disappear during the typical evolution of a system as it is heats up, evolves in time, or as its parts become separated. These results follow from the nature of the entanglement-free continent in the space of physical states, and hold in great generality. We illustrate these phenomena with a frustrated molecular quantum magnet in and out of equilibrium. In contrast, if the particles are fermions, such as electrons, another notion of entanglement exists that precludes entanglement-free regions, and thus protects quantum correlations. These findings provide fundamental knowledge about the structure of entanglement in quantum matter and architectures, paving the way for its manipulation.
Introduction.— Entanglement is a non-local quantum correlation between two or more particles that makes it possible for a measurement on a subset of particles to affect subsequent measurements on the other particles. The effect of the initial measurement is instantaneous, even if the particles are distant. Entanglement not only constitutes a fundamental property of nature, but it is also a resource to perform tasks that would prove impossible without it such as teleportation bennett1993teleporting , or more broadly quantum computation nielsen2010quantum . This has driven the community to devise methods to detect and quantify entanglement vedral1997quantifying . Unfortunately, it it not known how to determine with certainty whether a general system is entangled, except in very simple situations such as with 2 qubits. One can better grasp the complexity of the task by observing that entanglement can exist between more than two parties. In fact, some systems possess 3-party entanglement but no 2-party entanglement of any sort. In this work, we find criteria for such collective non-local quantum correlations in physical systems under very general conditions. For instance, at what temperatures can multi-party entanglement of a given kind exist? We begin by explaining important properties about the space of physical states, and how these determine the fate of entanglement under the evolution of a system with temperature, time or separation. We then illustrate these results with a simple yet generic model: the anti-ferromagnetic Ising model on an icosahedral molecule. Finally, we discuss the fate of entanglement for fermionic systems, where the parity superselection rule greatly modifies the geometry of the space of states and the structure of entanglement.
@@
where $\rho_{j}$ is a physical density matrix for subsystem $j$. The set of separable sates is convex vedral1997quantifying , namely states composed of a mixture of product states, $\rho_{\rm sep}=\sum_{k}p_{k}\rho_{\rm prod}^{(k)}$ with $p_{k}\geqslant 0$, are also un-entangled. In the space of all physical states, separable ones thus form a separable continent surrounded by an ocean of entangled states, see Fig. 1b. Highly-entangled states live in the deep-blue waters, whereas in the center of the separable continent lies the maximally-mixed state, i.e., the most classical state obtained in the limit of infinite temperature.
We describe the system’s state by $\rho(s)$, where $s$ parametrizes the evolution; the final state is $\rho_{f}\equiv\rho(\infty)$. We shall argue that in situations of physical relevance the final state is generically separable. Moreover, $\rho_{f}$ typically lies in the interior of the separable continent, not on its frontier. We thus arrive at our main conclusion: the system irreversibly looses all forms of multi-party entanglement beyond some stage in the evolution, as shown in Fig. 1b. Moreover, before reaching land, the system navigates shallow waters, leading to a rapid decimation of entanglement. During this stage one has an effectively separable state. This constitutes a point of significant importance in real-world applications since determining when a state reaches a very low degree of entanglement is easier than showing exact separability. Nevertheless, we can exploit a mathematical result that tells us when a state is in the interior of the separable continent braunstein1999separability ; gurvits2002largest ; shen2015bipartite ; wen2023separable . In the simplest version, the argument states that a product state (1) lies in the interior if it has full rank, i.e., none of its eigenvalues vanishes. This is coherent with what is known for pure states: these have minimal rank (a single non-zero eigenvalue), and indeed live on the frontier of the continent as arbitrarily weak perturbations make them entangled. Furthermore, the argument gives us the radius of a ball in the space of states that lies entirely on the separable continent; this is represented by the dashed line in Fig. 1b. The radius of the ball is proportional to the smallest eigenvalue of the state wen2023separable , $R_{m}=2^{1-m/2}\lambda_{\rm min}$. Here, we use the standard notion of distance between two quantum states given by the Frobenius norm: $d(\rho,\rho^{\prime})=\sqrt{\operatorname{Tr}(\rho-\rho^{\prime})^{2}}$.
@@
Temperature.— In the first application, we take the parameter $s$ to represent the temperature $T$. Naturally, a very large temperature destroys entanglement, yielding a maximally uncorrelated state. The infinite-temperature end point is a full-rank product state, $\rho_{f}=D^{-1}\mathbb{I}$, where $\mathbb{I}$ represents the identity matrix in the $D$-dimensional space of the system. This state is located at the center of the separable continent. Hence, there exists a threshold temperature $T_{m}$ at which we can rigorously conclude that all forms of entanglement between the $m$ subsystems disappear. The minimal eigenvalue of $\rho_{f}$ being $1/D$, one can readily obtain a temperature above which entanglement is lost. However, in the example below we shall see that the system rapidly enters shallow waters, and becomes effectively separable at much lower temperatures. Sudden death of bipartite entanglement at finite temperature has been observed in different physical systems arnesen2001natural ; anders2008thermal ; gong2009thermal ; hart2018entanglement ; ma2022symmetric .
@@
We can argue in full generality that the entanglement dynamics after a quench from a pure product state typically follows a rise-and-fall behavior. For early times, we consider the evolution of the system with parameter $s=t^{-1}$. The “final” state of this evolution is the initial state of the quench, namely a pure product state with rank one, i.e., not full rank. As we previously discussed, such states lie on the boundary, or the shore, of the separable continent. Therefore, there is no entanglement sudden death in this evolution. Looking back at the quench protocol with time as the parameter, this implies that entanglement is generated at $t=0^{+}$ after the quench. As discussed above, at late times the state typically lands on the continent, or in very shallow waters at a finite time, yielding an entanglement sudden death or strong suppression. These early- and late-time behaviors result in the rise-and-fall dynamics of entanglement. Such behavior has been observed for simple entanglement measures in numerous systems yu2009sudden ; ma2012entanglement ; alba2019quantum ; murciano2022quench ; PBC22 ; parez2023analytical ; sang2023ultrafast .
@@
Space.— We now study the fate of entanglement as a function of the separation between the $m$ subsystems by scaling their separations by $\lambda$, which parametrizes the evolution. At large $\lambda$, in a local physical system the state will factorise into a product form, where the description of each subsystem becomes independent, since all correlations become suppressed at large separations. If the asymptotic product state is of full rank, there exists a critical scale beyond which all entanglement vanishes. An important application is to study the entanglement between $m$ subsystems embedded in a larger system as these are chosen to be progressively further apart. The state at infinite $\lambda$ satisfies the product form, Eq. (1), where $\rho_{j}$ is the reduced density matrix of subsystem $j$ obtained by tracing out the complementary degrees of freedom. Let us take the entire system to be in equilibrium at a temperature $T~{}\!\!\geqslant~{}\!\!0$. At large separation, the state of subsystem $A_{1}$ is given by $\rho_{1}=\operatorname{Tr}_{A_{2}...A_{m}B}\sum_{n}p_{n}|E_{n}\rangle\langle E_{n}|$, where we have traced over the environment $B$, and the $(m-1)$ subsystems $A_{j}$. The sum runs over all eigenstates of the Hamiltonian. The thermal Boltzmann probabilities are $p_{n}=e^{-E_{n}/T}/\mathcal{Z}$. We now need to determine whether $\rho_{1}$ has full rank. Let us first examine the restriction of the eigenstates of the Hamiltonian to $A_{1}$, $\rho_{1}^{(n)}=\operatorname{Tr}_{A_{2}...A_{m}B}|E_{n}\rangle\langle E_{n}|$. From the point of view of $A_{1}$, the degrees of freedom in $B$ act as a bath, thus introducing statistical randomness in $\rho_{1}^{(n)}$. Since we take $A_{1}$ to be sufficiently small compared to the complement, the bath generically has sufficient resources to induce statistical fluctuations that span the entire Hilbert space of $A_{1}$. Such ergodicity implies that $\rho_{1}^{(n)}$ will be of full rank. This indeed occurs for the majority of eigenstates as they live near the middle of the energy spectrum. The eigenstate thermalization hypothesis deutsch1991quantum ; srednicki1994chaos ; rigol2008thermalization then states that the reduced density matrix on $A_{1}$ is approximately thermal, with the temperature determined by the energy $E_{n}$. The approximately thermal density matrix, obtained by restricting the Hamiltonian to $A_{1}$, then has full rank. Our statement, which we call the full-rank hypothesis (FRH), is more general: all eigenstates of a generic local Hamiltonian have a full-rank reduced density matrix associated with a sufficiently small subregion. In practice, the subregion should not exceed half the system.
@@
Now, the reduced density matrix of subsystem 1 is the convex sum $\rho_{1}=\sum_{n}p_{n}\rho_{1}^{(n)}$. The FRH implies that $\rho_{1}$ inherits full rank from the states $\rho_{1}^{(n)}$. (Since the sum of a full-rank matrix and an arbitrary one also has full rank, we actually only need to know that at least one $\rho_{1}^{(n)}$ is of full rank.) As the same holds for the other subsystems, we conclude that the $\lambda=\infty$ product has full rank, and thus lies on the separable continent. Therefore, there exists a scale beyond which all entanglement disappears. The sudden death occurs irrespective of the number of subsystems or the precise nature of the state; it even holds at a quantum critical phase transition where quantum fluctuations proliferate to all scales. This result shows that, although local observables can possess slowly decaying algebraic correlations, entanglement decays drastically faster, to the point of having a finite range. Our conclusion encompasses and generalises numerous examples of bipartite entanglement sudden death at finite separation Osterloh_2002 ; Osborne ; JTBBJH18 ; klco2021entanglement ; klco2021geometric ; parez2023separability ; parez2023fermionic , which are very specific cases of the general multi-partite phenomenon.
@@
Icosahedral molecule.— We illustrate the above results with a simple quantum system: the anti-ferromagnetic Ising model on the 12-spin icosahedron, see Fig. 2a, with Hamiltonian $H=J\sum_{{\rm bonds}\,\langle i,j\rangle}\sigma_{i}^{x}\sigma_{j}^{x}-h\sum_{{\rm sites}\,i}\sigma_{i}^{z}$. The first term corresponds to an anti-ferromagnetic interaction $J>0$ that tends to anti-align neighbouring spins along $x$, while the second one is a transverse field that polarises all spins along $z$. Such a molecular quantum magnet is a combination of identical triangular faces, and thus possesses strong geometric frustration. In what follows we shall measure energy in units of the exchange coupling by setting $J=1$. We investigate the fate of 2- and 3-spin entanglement as a function of magnetic field, temperature, time a separation by considering three distinct quantities. See the Supplemental Material (SM) for their precise definitions. First, we compute a measure of 2-party entanglement, focusing on a pair of adjacent spins, called the logarithmic negativity $\mathcal{E}$ VW02 ; plenio2005logarithmic . In the case of two spins, $\mathcal{E}=0$ implies that the reduced density matrix is separable HORODECKI19961 , whereas for entangled states we have $\mathcal{E}>0$. Second, we study the 3-spin entanglement on a triangular face of the molecule via the geometric entanglement $\mathcal{D}$ vedral1997quantifying defined as the smallest distance between the state and the separable continent,
@@
where the minimisation is over all $\rho_{\rm sep}$ living on the continent. This powerful measure detects all forms of entanglement. Finally, we employ a quantity $W$ that indicates the presence of genuine 3-party entanglement when $W>0$ guhne2010separability . This type of entanglement relates all three parties, and cannot be described in term 2-party entanglement. When $W=0$, a definitive conclusion cannot be made. We note that both $\mathcal{D}$ and $W$ require a minimisation procedure in a multi-parameter space, making their evaluation more demanding than simpler measures such as $\mathcal{E}$. As a prerequisite, we checked that the FRH holds for subregions of adjacent spins with 1, 2, 3, and 4 sites by obtaining the $2^{12}$ eigenstates via exact numerical diagonalization.
@@
Let us now investigate the fate of entanglement with temperature at $h=3$. From the criterion of the separable ball around the infinite-temperature state wen2023separable , we determine the temperatures where $m$-party entanglement is guaranteed to be absent for $m=2,3$. We find $T_{2}=8$ and $T_{3}=22$, respectively. However, these do not tell us if entanglement is absent at lower temperatures. The logarithmic negativity $\mathcal{E}$ and the witness $W$ vanish a temperatures $T_{\mathcal{E}}=2.1$ and $T_{W}=1.5$, which are much smaller than $T_{2}$ and $T_{3}$. Finally, we observe that for $T_{3}^{*}\approx 3$, the geometric entanglement $\mathcal{D}$ is of order $10^{-4}$, and further rapidly decreases with temperature beyond that point. Hence, the 3-spin density matrix is essentially separable for temperatures much smaller than the rigorous bound $T_{3}$. We report these results in Fig. 2c.
Next, we consider the evolution of entanglement with time during a quantum quench. We initialise the system in a pure product state of up and down spins in the $\sigma^{z}$-basis at $t=0$, and let it evolve under the Ising Hamiltonian discussed above with $h=3$. In Fig. 2d we show the evolution of our three measures of interest. Both $\mathcal{E}$ and $W$ rapidly rise, reach a maximum, and fall to zero, with $W$ being the shortest lived. We thus see a clear illustration of the sudden death of entanglement for the case of $m=2$ spins. The distance $\mathcal{D}$ follows the same behavior but displays small oscillations close to zero for later times. The state quickly reaches the shallow waters around $t_{3}^{*}\approx 0.4$, while an exact sudden death does not occur for $m=3$, presumably due to the small size of the molecule. An interesting observation is that the non-equilibrium evolution can generate stronger 2- and 3-party entanglement compared to what is found in equilibrium, even at zero temperature, see Fig. 2c. Finally, we observe qualitatively identical behavior for other initial conditions.
@@
Figure 2: Icosahedral molecule and its entanglement evolutions. a) Illustration of the quantum molecular magnet with icosahedral geometry. b) Dependence of the geometric entanglement $\mathcal{D}$, the 2-spin logarithmic negativity $\mathcal{E}$ and the genuine 3-party entanglement witness $W$ in the icosahedral molecule as a function of the transverse field at zero temperature. c) Evolution of the same quantities as a function of temperature with $h=3$. d) Same quantities as a function of time in a quench protocol where the system is prepared in a product state and evolves under the anti-ferromagnetic Ising Hamiltonian with $h=3$.
@@
No continent for fermions.— The above discussion holds for systems of quantum spins. Crucially, spin operators at different sites are independent (they commute with each other) because they represent bosons. However, there exist other particles in nature that do not commute, instead they acquire a minus sign upon exchange: fermions, like electrons or protons. Individual fermion operators possess this relative non-locality with other fermions, but observers nevertheless witness a local world since physical operators are made of an even number of fermions, and are thus bosonic. The density matrix of a physical fermionic system made of $m$ subsystems thus possesses even fermion parity. But how does fermion parity affects the definition of entanglement by constraining viable separable states? A natural notion follows from declaring that a fermionic $m$-party state is un-entangled if it can be written as $\rho^{F}_{\rm sep}=\sum_{k}p_{k}\rho^{(k)}_{1}\otimes\rho^{(k)}_{2}\otimes\cdots\otimes\rho^{(k)}_{m}$, where each $\rho^{(k)}_{j}$ has even parity and the $p_{k}$ form a probability distribution banuls2007entanglement ; spee2018mode ; shapourian2019entanglement . This definition guarantees that the state has no quantum correlations among the $m$ components, and can thus be prepared locally. Certain states can be brought to this form, but where some $\rho^{(k)}_{j}$do not have even parity. These states cannot be prepared locally and their non-local correlations can be exploited for quantum tasks such as quantum teleportation or quantum data hiding verstraete2003quantum ; schuch2004nonlocal ; schuch2004quantum .
@@
Given an un-entangled fermionic physical state $\rho^{F}_{\rm sep}$, can we find an entanglement-free region around it as was the case for bosons? We show that the answer is no, and then discuss the consequences. The argument, as given in Ref. benatti2014entanglement , is the following. It suffices to consider the case with $m=2$ components by performing a small deformation, $\rho^{F}_{\rm sep}+\delta\rho$, where $\delta\rho$ has zero trace and contains terms of odd fermion parity for subsystem 1. For example, one can hop an odd number of fermions between subsystems 1 and 2, e.g., $\delta\rho=\epsilon(|0_{1}\rangle\langle 1_{2}|+|1_{1}\rangle\langle 0_{2}|)$ with positive $\epsilon\ll 1$. The deformed state does not have even parity for subsystem 1, and hence is not separable, even for arbitrarily small $\epsilon$. Therefore, no entanglement-free region exists around a fermionic separable state since nearby states arbitrarily close to $\rho^{F}_{\rm sep}$ contain entanglement between any two subsystems. In the fermionic world, there is thus no separable continent. Separable states lie on a zero-width sand beach, surrounded by an ocean of entangled states arbitrarily close by. We illustrate this in Fig 3.
@@
Turning the tables around, fermionic entanglement is more robust since the state cannot land on a continent upon evolution. Let us say that we heat a fermion system that contains multi-party entanglement at low $T$. The infinite temperature state is un-entangled, but entanglement will decay gradually as a function of temperature without a sudden death. This behavior was observed for integer quantum Hall states liu2022entanglement and free-fermion systems banuls2007entanglement ; choi2023finite . Analogous conclusions hold for temporal or spatial evolution. For instance, the entanglement between disjoint regions in fermionic conformal field theories in arbitrary dimensions decays as a power-law with the distance, without a sudden death parez2023fermionic .
@@
Outlook.— In this paper, we discussed how the structure of the space of physical states, and in particular the presence of a separable continent, leads to the sharp disappearance of entanglement in various physical situations, under very generic assumptions. For fermionic systems, the parity superselection rule forbids the existence of the separable continent and drastically modifies the fate of entanglement. In particular, fermionic systems do not experience entanglement sudden death, in stark contrast with bosonic ones. Interestingly, the same conclusion holds for other types of superselection rules, providing a generic recipe for the creation of robust non-local entangled states.

### v2

@@ Abstract
Quantum entanglement is a fundamentally non-local correlation between particles. In its simplest realisation, a measurement on one particle is affected by a prior measurement on its partner, irrespective of their separation. For multiple particles, purely collective types of entanglement exist but their detection, even theoretically, remains an outstanding open question. Here, we show that all forms of multi-party entanglement entirely disappear during the typical evolution of a system as it heats up, evolves in time, or as its parts become separated. These results follow from the nature of the entanglement-free continent in the space of physical states, and hold in great generality. We illustrate these phenomena with a frustrated molecular quantum magnet in and out of equilibrium. In contrast, if the particles are fermions, such as electrons, another notion of entanglement exists that precludes entanglement-free regions, and thus protects quantum correlations. These findings provide fundamental knowledge about the structure of entanglement in quantum matter and architectures, paving the way for its manipulation.
Introduction.— Entanglement is a non-local quantum correlation between two or more particles that makes it possible for a measurement on a subset of particles to affect subsequent measurements on the other particles. The effect of the initial measurement is instantaneous, even if the particles are distant. Entanglement not only constitutes a fundamental property of nature, but it is also a resource to perform tasks that would prove impossible without it such as teleportation [1], or more broadly quantum computation [2]. This has driven the community to devise methods to detect and quantify entanglement [3]. Unfortunately, it it not known how to determine with certainty whether a general system is entangled, except in very simple situations such as with 2 qubits. One can better grasp the complexity of the task by observing that entanglement can exist between more than two parties. In fact, some systems possess 3-party entanglement but no 2-party entanglement of any sort. In this work, we find criteria for such collective non-local quantum correlations in physical systems under very general conditions. For instance, at what temperatures can multi-party entanglement of a given kind exist? We begin by explaining important properties about the space of physical states, and how these determine the fate of entanglement under the evolution of a system with temperature, time or separation. We then illustrate these results with a simple yet generic model: the anti-ferromagnetic Ising model on an icosahedral molecule. Finally, we discuss the fate of entanglement for fermionic systems, where the parity superselection rule greatly modifies the geometry of the space of states and the structure of entanglement.
@@
where $\rho_{j}$ is a physical density matrix for subsystem $j$. The set of separable sates is convex [3], namely states composed of a mixture of product states, $\rho_{\rm sep}=\sum_{k}p_{k}\rho_{\rm prod}^{(k)}$ with $p_{k}\geqslant 0$, are also un-entangled. In the space of all physical states, separable ones thus form a separable continent surrounded by an ocean of entangled states, see Fig. 1b. Highly-entangled states live in the deep-blue waters, whereas in the center of the separable continent lies the maximally-mixed state, i.e., the most classical state obtained in the limit of infinite temperature.
We describe the system’s state by $\rho(s)$, where $s$ parametrizes the evolution; the final state is $\rho_{f}\equiv\rho(\infty)$. We shall argue that in situations of physical relevance the final state is generically separable. Moreover, $\rho_{f}$ typically lies in the interior of the separable continent, not on its frontier. We thus arrive at our main conclusion: the system irreversibly looses all forms of multi-party entanglement beyond some stage in the evolution, as shown in Fig. 1b. Moreover, before reaching land, the system navigates shallow waters, leading to a rapid decimation of entanglement. During this stage one has an effectively separable state. This constitutes a point of significant importance in real-world applications since determining when a state reaches a very low degree of entanglement is easier than showing exact separability. Nevertheless, we can exploit a mathematical result that tells us when a state is in the interior of the separable continent [4, 5, 6, 7]. In the simplest version, the theorem states that a product state (1) lies in the interior if it has full rank, i.e., none of its eigenvalues vanishes. This is coherent with what is known for pure product states: these have minimal rank (a single non-zero eigenvalue), and indeed live on the frontier of the continent as arbitrarily weak perturbations can make them entangled. Furthermore, the argument gives us the radius of a ball in the space of states that lies entirely on the separable continent; this is represented by the dashed line in Fig. 1b. The radius of the ball is proportional to the smallest eigenvalue of the state [7], $R_{m}=2^{1-m/2}\lambda_{\rm min}$. Here, we use the standard notion of distance between two quantum states given by the Frobenius norm: $d(\rho,\rho^{\prime})=\sqrt{\operatorname{Tr}(\rho-\rho^{\prime})^{2}}$. A more general version of the result holds for a mixture of product states with a least one having full rank [7].
@@
Temperature.— In the first application, we take the parameter $s$ to represent the temperature $T$. Naturally, a very large temperature destroys entanglement, yielding a maximally uncorrelated state. The infinite-temperature end point is a full-rank product state, $\rho_{f}=D^{-1}\mathbb{I}$, where $\mathbb{I}$ represents the identity matrix in the $D$-dimensional space of the system. This state is located at the center of the separable continent. Hence, there exists a threshold temperature $T_{m}$ at which we can rigorously conclude that all forms of entanglement between the $m$ subsystems disappear. The minimal eigenvalue of $\rho_{f}$ being $1/D$, one can readily obtain a temperature above which entanglement is lost. However, in the example below we shall see that the system rapidly enters shallow waters, and becomes effectively separable at much lower temperatures. Sudden death of bipartite entanglement at finite temperature has been observed in different physical systems [8, 9, 10, 11, 12].
@@
We can argue in full generality that the entanglement dynamics after a quench from a pure product state typically follows a rise-and-fall behavior. For early times, we consider the evolution of the system with parameter $s=t^{-1}$. The “final” state of this evolution is the initial state of the quench, namely a pure product state with rank one, i.e., not full rank. As we previously discussed, such states lie on the boundary, or the shore, of the separable continent. Therefore, there is no entanglement sudden death in this evolution. Looking back at the quench protocol with time as the parameter, this implies that entanglement is generated at $t=0^{+}$ after the quench. As discussed above, at late times the state typically lands on the continent, or in very shallow waters at a finite time, yielding an entanglement sudden death or strong suppression. These early- and late-time behaviors result in the rise-and-fall dynamics of entanglement. Such behavior has been observed for simple entanglement measures in numerous systems [13, 14, 15, 16, 17, 18, 19].
@@
Space.— We now study the fate of entanglement as a function of the separation between the $m$ subsystems by scaling their separations by $\lambda$, which parametrizes the evolution. At large $\lambda$, in a local physical system the state will factorise into a product form, where the description of each subsystem becomes independent, since all correlations become suppressed at large separations. If the asymptotic product state is of full rank, there exists a critical scale beyond which all entanglement vanishes. An important application is to study the entanglement between $m$ subsystems embedded in a larger system as these are chosen to be progressively further apart. The state at infinite $\lambda$ satisfies the product form, Eq. (1), where $\rho_{j}$ is the reduced density matrix of subsystem $j$ obtained by tracing out the complementary degrees of freedom. Let us take the entire system to be in equilibrium at a temperature $T~{}\!\!\geqslant~{}\!\!0$. At large separation, the state of subsystem $A_{1}$ is given by $\rho_{1}=\operatorname{Tr}_{A_{2}...A_{m}B}\sum_{n}p_{n}|E_{n}\rangle\langle E_{n}|$, where we have traced over the environment $B$, and the $(m-1)$ subsystems $A_{j}$. The sum runs over all eigenstates of the Hamiltonian. The thermal Boltzmann probabilities are $p_{n}=e^{-E_{n}/T}/\mathcal{Z}$. We now need to determine whether $\rho_{1}$ has full rank. Let us first examine the restriction of the eigenstates of the Hamiltonian to $A_{1}$, $\rho_{1}^{(n)}=\operatorname{Tr}_{A_{2}...A_{m}B}|E_{n}\rangle\langle E_{n}|$. From the point of view of $A_{1}$, the degrees of freedom in $B$ act as a bath, thus introducing statistical randomness in $\rho_{1}^{(n)}$. Since we take $A_{1}$ to be sufficiently small compared to the complement, the bath generically has sufficient resources to induce statistical fluctuations that span the entire Hilbert space of $A_{1}$. Such ergodicity implies that $\rho_{1}^{(n)}$ will be of full rank. This indeed occurs for the majority of eigenstates as they live near the middle of the energy spectrum. The eigenstate thermalization hypothesis [20, 21, 22] then states that the reduced density matrix on $A_{1}$ is approximately thermal, with the temperature determined by the energy $E_{n}$. The approximately thermal density matrix, obtained by restricting the Hamiltonian to $A_{1}$, then has full rank. Our statement, which we call the full-rank hypothesis (FRH), is more general: all eigenstates of a generic local Hamiltonian have a full-rank reduced density matrix associated with a sufficiently small subregion. In practice, the subregion should not exceed half the system.
@@
Now, the reduced density matrix of subsystem 1 is the convex sum $\rho_{1}=\sum_{n}p_{n}\rho_{1}^{(n)}$. The FRH implies that $\rho_{1}$ inherits full rank from the states $\rho_{1}^{(n)}$. (Since the sum of a full-rank matrix and an arbitrary one also has full rank, we actually only need to know that at least one $\rho_{1}^{(n)}$ is of full rank.) As the same holds for the other subsystems, we conclude that the $\lambda=\infty$ product has full rank, and thus lies on the separable continent. Therefore, there exists a scale beyond which all entanglement disappears. The sudden death occurs irrespective of the number of subsystems or the precise nature of the state; it even holds at a quantum critical phase transition where quantum fluctuations proliferate to all scales. This result shows that, although local observables can possess slowly decaying algebraic correlations, entanglement decays drastically faster, to the point of having a finite range. Our conclusion encompasses and generalises numerous examples of bipartite entanglement sudden death at finite separation [23, 24, 25, 26, 27, 28, 29], which are very specific cases of the general multi-party phenomenon.
@@
Icosahedral molecule.— We illustrate the above results with a simple quantum system: the anti-ferromagnetic Ising model on the 12-spin icosahedron, see Fig. 2a, with Hamiltonian $H=J\sum_{{\rm bonds}\,\langle i,j\rangle}\sigma_{i}^{x}\sigma_{j}^{x}-h\sum_{{\rm sites}\,i}\sigma_{i}^{z}$. The first term corresponds to an anti-ferromagnetic interaction $J>0$ that tends to anti-align neighbouring spins along $x$, while the second one is a transverse field that polarises all spins along $z$. Such a molecular quantum magnet is a combination of identical triangular faces, and thus possesses strong geometric frustration. In what follows we shall measure energy in units of the exchange coupling by setting $J=1$. We investigate the fate of 2- and 3-spin entanglement as a function of magnetic field, temperature, time a separation by considering three distinct quantities. See the Supplemental Material (SM) for their precise definitions. First, we compute a measure of 2-party entanglement, focusing on a pair of adjacent spins, called the logarithmic negativity $\mathcal{E}$ [30, 31]. In the case of two spins, $\mathcal{E}=0$ implies that the reduced density matrix is separable [32], whereas for entangled states we have $\mathcal{E}>0$. Second, we study the 3-spin entanglement on a triangular face of the molecule via the geometric entanglement $\mathcal{D}$ [3] defined as the smallest distance between the state and the separable continent,
@@
where the minimisation is over all $\rho_{\rm sep}$ living on the continent. This powerful measure detects all forms of entanglement. Finally, we employ a criterion $W$ that indicates the presence of genuine 3-party entanglement when $W>0$ [33]. This type of entanglement relates all three parties, and cannot be described in term 2-party entanglement. When $W=0$, a definitive conclusion cannot be made. We note that both $\mathcal{D}$ and $W$ require a minimisation procedure in a multi-parameter space, making their evaluation more demanding than simpler measures such as $\mathcal{E}$. As a prerequisite, we checked that the FRH holds for subregions of adjacent spins with 1, 2 and 3 sites by obtaining the $2^{12}$ eigenstates via exact numerical diagonalization, see SM.
@@
Let us now investigate the fate of entanglement with temperature at $h=3$. From the criterion of the separable ball around the infinite-temperature state [7], we determine the temperatures where $m$-party entanglement is guaranteed to be absent for $m=2,3$. We find $T_{2}=8$ and $T_{3}=22$, respectively. However, these do not tell us if entanglement is absent at lower temperatures. The logarithmic negativity $\mathcal{E}$ and the criterion $W$ vanish at temperatures $T_{\mathcal{E}}=2.1$ and $T_{W}=1.5$, which are much smaller than $T_{2}$ and $T_{3}$. Finally, we observe that for $T_{3}^{*}\approx 3$, the geometric entanglement $\mathcal{D}$ is of order $10^{-4}$, and further rapidly decreases with temperature beyond that point. Hence, the 3-spin density matrix is essentially separable for temperatures much smaller than the rigorous bound $T_{3}$. We report these results in Fig. 2c.
Next, we consider the evolution of entanglement with time during a quantum quench. We initialise the system in a pure product state of up and down spins in the $\sigma^{z}$-basis at $t=0$, and let it evolve under the Ising Hamiltonian discussed above with $h=3$, see SM for details regarding the initial state. In Fig. 2d we show the evolution of our three measures of interest. Both $\mathcal{E}$ and $W$ rapidly rise, reach a maximum, and fall to zero, with $W$ being the shortest lived. We thus see a clear illustration of the sudden death of entanglement for the case of $m=2$ spins. The distance $\mathcal{D}$ follows the same behavior but displays small oscillations close to zero for later times. The state quickly reaches the shallow waters around $t_{3}^{*}\approx 0.4$, while an exact sudden death does not occur for $m=3$, presumably due to the small size of the molecule. An interesting observation is that the non-equilibrium evolution can generate stronger 2- and 3-party entanglement compared to what is found in equilibrium, even at zero temperature, see Fig. 2c. Finally, we observe qualitatively identical behavior for other initial conditions.
@@
Figure 2: Icosahedral molecule and its entanglement evolutions. a) Illustration of the quantum molecular magnet with icosahedral geometry. b) Dependence of the geometric entanglement $\mathcal{D}$, the 2-spin logarithmic negativity $\mathcal{E}$ and the genuine 3-party entanglement criterion $W$ in the icosahedral molecule as a function of the transverse field at zero temperature. c) Evolution of the same quantities as a function of temperature with $h=3$. d) Same quantities as a function of time in a quench protocol where the system is prepared in a product state and evolves under the anti-ferromagnetic Ising Hamiltonian with $h=3$.
@@
No continent for fermions.— The above discussion holds for systems of quantum spins. Crucially, spin operators at different sites are independent (they commute with each other) because they represent bosons. However, there exist other particles in nature that do not commute, instead they acquire a minus sign upon exchange: fermions, like electrons or protons. Individual fermion operators possess this relative non-locality with other fermions, but observers nevertheless witness a local world since physical operators are made of an even number of fermions, and are thus bosonic. The density matrix of a physical fermionic system made of $m$ subsystems thus possesses even fermion parity. But how does fermion parity affects the definition of entanglement by constraining viable separable states? A natural notion follows from declaring that a fermionic $m$-party state is un-entangled if it can be written as $\rho^{F}_{\rm sep}=\sum_{k}p_{k}\rho^{(k)}_{1}\otimes\rho^{(k)}_{2}\otimes\cdots\otimes\rho^{(k)}_{m}$, where each $\rho^{(k)}_{j}$ has even parity and the $p_{k}$ form a probability distribution [34, 35, 36]. This definition guarantees that the state has no quantum correlations among the $m$ components, and can thus be prepared locally. Certain states can be brought to this form, but where some $\rho^{(k)}_{j}$do not have even parity. These states cannot be prepared locally and their non-local correlations can be exploited for quantum tasks such as quantum teleportation or quantum data hiding [37, 38, 39].
@@
Given an un-entangled fermionic physical state $\rho^{F}_{\rm sep}$, can we find an entanglement-free region around it as was the case for bosons? We show that the answer is no, and then discuss the consequences. The argument, as given in Ref. [40], is the following. It suffices to consider the case with $m=2$ components by performing a small deformation, $\rho^{F}_{\rm sep}+\delta\rho$, where $\delta\rho$ has zero trace and contains terms of odd fermion parity for subsystem 1. For example, one can hop an odd number of fermions between subsystems 1 and 2, e.g., $\delta\rho=\epsilon(|01\rangle\langle 10|+|10\rangle\langle 01|)$ with positive $\epsilon\ll 1$. The deformed state does not have even parity for subsystem 1, and hence is not separable, even for arbitrarily small $\epsilon$. Therefore, no entanglement-free region exists around a fermionic separable state since nearby states arbitrarily close to $\rho^{F}_{\rm sep}$ contain entanglement between any two subsystems. In the fermionic world, there is thus no separable continent. Separable states lie on a zero-width sand beach, surrounded by an ocean of entangled states arbitrarily close by. We illustrate this in Fig 3.
@@
Turning the tables around, fermionic entanglement is more robust since the state cannot land on a continent upon evolution. Let us say that we heat a fermion system that contains multi-party entanglement at low $T$. The infinite temperature state is un-entangled, but entanglement will decay gradually as a function of temperature without a sudden death. This behavior was observed for integer quantum Hall states [41] and free-fermion systems [34, 42]. Analogous conclusions hold for temporal or spatial evolution. For instance, the entanglement between disjoint regions in fermionic conformal field theories in arbitrary dimensions decays as a power-law with the distance, without a sudden death [29].
@@
Outlook.— In this paper, we discussed how the structure of the space of physical states, and in particular the presence of a separable continent, leads to the sharp disappearance of entanglement in various physical situations, under very generic assumptions. For fermionic systems, the parity superselection rule forbids the existence of the separable continent and drastically modifies the fate of entanglement. In particular, fermionic systems do not experience entanglement sudden death, in stark contrast with bosonic ones. Interestingly, the same conclusion holds for other types of superselection rules, providing a generic recipe for the creation of robust entangled states.